OpenStudy (anonymous):

Establish the identity tan u cot u - cos^2u=sin^2u

hartnn (hartnn):

can you express tan u in terms of sin u and cos u ?

OpenStudy (anonymous):

Not really, I barely learned this today with a sub who couldn't teach..

hartnn (hartnn):

learn the basic identities first then? do you know any trigonometric identity ?

hartnn (hartnn):

for this Question, you'll need these, \(\huge \tan x=\dfrac{\sin x}{\cos x} \\ \huge \cot x=\dfrac{\cos x}{\sin x} \\ \huge \sin^2x+\cos^2x=1\)

OpenStudy (anonymous):

So.. sinx/cosx * cosx/sinx - cos^2u=sin^2u?

hartnn (hartnn):

yes, what gets cancelled out ?

OpenStudy (anonymous):

sinx/cosx*cosx/sinx

hartnn (hartnn):

what remains ?

OpenStudy (anonymous):

-cos^2u=sin^2u

hartnn (hartnn):

a/a =1 and nor 0, so after cancelling out, the '1' remains, so you have 1-cos^2 u on left, got this ?

OpenStudy (anonymous):

no... :(

hartnn (hartnn):

why ? what part is confusing you ?

OpenStudy (anonymous):

a/a =1 and nor 0, so after cancelling out, the '1' remains, ^ this part

hartnn (hartnn):

whats 2/2 =... ?

OpenStudy (anonymous):

1

hartnn (hartnn):

yes, infact anything/same thing = 1 so,  now got that^ ?